The next number in the pattern – 62, – 37, – 12 _________ is (a) 25 (b) 13 (c) 0 (d) –13

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student


Option (a) is correct here

It’s found that the pattern is-62 + 25 = -37, -37 + 25
= -12
So, similarly -12 + 25 = 13


Thanks